# Textoshop: Interactions Inspired by Drawing Software to Facilitate Text Editing

### v1

@@ 3.6.2. Prompting Strategies and Heuristics
Our intent with the implementation was to have robust, fast, and reliable tools to allow testing the concept behind Textoshop. While the implementation could perhaps be improved, we currently rely on engineered prompts and heuristics. For example, tools operating on a selection were optimized to only send the selection to the model and reintegrate the results by making it match the formatting of the original text. Another example is the resizing tool which first splits the selection into sentences and prompts the model to shorten or expand each sentence in parallel. In our tests, GPT-4o was unreliable when prompted to resize a text by a certain amount of words. To boost the reliability of the resizing, the model was prompted to resize the same text eight times using a slightly different number of words. Then, the system calculated all possible combinations of the resized parts and selected the one closest to the target length.
@@ 4. User Study
We ran a two-part user study to learn about the usability and usefulness of our approach. Specifically, we tried to answer the following questions: Can people use the features of Textoshop to accomplish typical writing tasks? Does their knowledge of drawing software help them understand how to use Textoshop? Is Textoshop easier to use than existing solutions?
@@
The second part of the study was more exploratory, allowing us to observe how participants use the tool as a whole in a more holistic and unguided way. Participants used the fully-featured version of Textoshop on their own text with no other instruction than to explore the tool and its features to accomplish their text editing goals.
@@ 4.1. Participants
We recruited 12 participants from our local community (age range: 23 to 35, Mdn=26; 7 self-identified as female and 5 as male). On a 5-point scale from 1-never to 5-very frequently, participants reported how often they engaged in writing activities (Mdn=5, SD=.7), how often they use generative AI to help them during writing activities (Mdn=2.5, SD=1.5), and how often they use drawing software (Mdn=3, SD=1). They indicated writing content related to academic writing (12), professional writing (11), and creative writing (2). On a 5-point scale from 1-not at all familiar to 5-extremely familiar, they also reported their familiarity with specific drawing software features such as colour pickers (Mdn=5, SD=.8), drawing tools (Mdn=4, SD=.7), layers (Mdn=4, SD=1.1), and boolean operations on shapes (Mdn=4, SD=.6).
@@ 4.3. Procedure
Each participant used two interfaces (Textoshop and baseline). For each interface, the participant completed 4 tasks (shortening and expanding a story, organizing and integrating text fragments, changing the tone of a dialogue, and using a template to write different emails).
Introduction (~2min) – At the start of the session, the participant completed the consent form and demographics questionnaire and was told they would use two different interfaces to accomplish a series of text editing tasks.
@@ 4.4. Design and Tasks
We designed four tasks such that they cover the issues with traditional ways of editing text (section 2.1) and involve as many of the features of Textoshop as possible. The tasks were designed to be realistic, require multiple steps, and ask of participants to express higher-level intents rather than simple editing operations. Rather than evaluating all the individual tools, our main question was whether participants understand and successfully use features building on the drawing software analogy. The 4 tasks were as follows.
@@
Shortening/Expanding a Summary – Given a summary of a well-known novel, the participant has to shorten and expand specific sentences and paragraphs. For example, they need to compress a paragraph to avoid the short last line (“orphan”) while preserving the meaning. Or they need to expand upon a specific sentence in another paragraph so that it covers two lines and fills the whole width of the page. With Textoshop, the participant can use direct manipulation to resize text selections. This task was to test precise resizing through direct manipulation.
@@
Changing the Tone of Dialogues – Given a dialogue between two fictional characters, the participant has to change the tone of the dialogue of one of these characters (e.g., make their dialogues more informal and positive) and then explore at least four variations of the other character’s last dialogue. With Textoshop, the participant can use the tone picker and tone-brush. This task was to test the tone picker metaphor to explore and find variations of tone.
@@
Organizing and Integrating Text Fragments – Given a simple story and a bullet list with 6 text fragments, the participant has to sort the text fragments depending on the character they are related to and then integrate these fragments into the story. Some text fragments have to be merged into specific parts of the text, while others are concepts to remove from the story. In the Textoshop condition, participants can arrange the text fragments in and off the page and can also use boolean operations to unite, intersect, subtract, or exclude text fragments. This task was to test the metaphor of the canvas and floating elements, as well as boolean operations on text fragments.
@@
Using an Email Template to Answer Customers – Given the profile of two customers as well as the email template used by a fictional company to answer their customers, the participant has to generate an email in response to each customer. The template includes pre-written paragraphs, which should be selected if they apply to that specific customer, and has to be edited further to include the name of the customer and some specific information, such as amounts and durations. With Textoshop, layers are already created with the text of the email template; to form the final emails, the participant has to show/hide different layers and apply minor edits to the layers. With the baseline interface, the email template lists all the possible phrasings to answer a customer organized in sections; to form the final emails, the participant has to copy and paste them in order. This task was to test the layer metaphor and participants’ mental model of layers when editing.
@@ 4.5. Data Analysis
Considering ongoing debates about statistics in HCI (Dragicevic, 2016; Group, 2019), we report both confidence intervals (CI) and p-values. However, we base our conclusions on the 95% confidence intervals (CI) on mean differences, which provide information about both the presence and size of effects. All confidence intervals are calculated using the studentized bootstrapping method because it is the most robust for study designs similar to ours (Masson et al., 2023; Zhu and Kolassa, 2018). All p-values are calculated using a Wilcoxon signed-rank test because it is less likely to yield false positives compared to parametric alternatives when dealing with data whose distribution is unknown and possibly heavily skewed (Bridge and Sawilowsky, 1999). The bootstrapped CIs are calculated in Python using arch 5.5.0 (Sheppard et al., 2023) and using 10,000 replications. The Wilcoxon signed-rank tests are calculated using scipy 1.10.1 (Virtanen et al., 2020) with ties being discarded.
@@ 4.6. Results from the Text Editing Tasks
Stacked bar chart of participants responses to 5-point statements with the mean difference showing on the right. Overall, ratings for Textoshop are mostly positive while those for baseline are mixed or negative.
@@ 4.6.4. Preferences
When asked about their preferred interface for each task, participants unanimously rated Textoshop as “preferred” or “strongly preferred” when “integrating text fragments”. All participants also preferred Textoshop when “precisely shortening/expanding text” except P10, who did not have a preference for either interface. For other tasks, while the majority preferred Textoshop, some participants expressed other preferences: when “writing using a template” P8 preferred the baseline interface; when “organizing text fragments” P3 preferred and P8 strongly preferred the baseline; and when “exploring meaningful variations of text” P7, P8, and P12 preferred and P3 strongly preferred the baseline. We discuss the reasons invoked by the participants for these preferences in Section 5.
@@ 4.7. Results from the Free-form Exploration
In the free-form exploration, participants chose texts that could be classified as creative story writing (P2, P4, P5, P9), informative writing (P3, P10, P11, P12), academic writing (P1, P6, P7), and personal writing (P8). Below, we report some of the strategies and behaviours we observed.
@@ 4.7.1. Iterating on a sentence by alternating tools
All participants used the tools to iterate on the wording of some specific sentences. Often, their workflow involved alternating between trying different tones, shortening the sentence, and smudging some parts to be paraphrased. The order of these steps was not always the same, but the tone picker was often used first, and then the other tools would allow editing with more precision. With the exception of P2, who frequently rotated and split the sentences, other tools were seldom used to rewrite a sentence.
@@ 4.7.2. Experimenting with alternative organizations of ideas
P5, P10, and P12 reorganized the ideas multiple times by taking a passage from one paragraph and integrating it into a different paragraph or sentence using the “unite” tool. P6 did the same but created their own text fragments to add an idea to an existing paragraph. Finally, P1 went one step further and deconstructed a whole paragraph into passages. Then, they organized these passages in the margin to try a different organization of the arguments and then reintegrated the passages into the main text using the “insert” and “unite” features.
@@ 4.7.3. Using layers to identify changes, store AI modifications, or add structure
Layers were used in different ways. For example, P1, P8, and P9 used one layer to store all the modifications made using the tools, allowing them to preserve the integrity of their original text and easily identify AI changes by switching the visibility of the layer. P3 used layers to store alternative versions of paragraphs (in this case, storing a more informal version). Finally, other participants used layers to structure their text. For example, P5, P11, and P12 created layers to store the different paragraphs, passages, or main ideas. P7 used layers to store the examples mentioned in the text. And P6 used a layer to store the statistics reported in the text so that they could read and edit their results section while hiding the means or p-values reported in the text (by hiding the statistics layer).
@@ 5. Discussion
We first summarize our study results with participants’ comments to answer our main research questions. Then, we open the discussion to other themes stemming from observations and discussions with participants.
Participants could use Textoshop to accomplish writing tasks
This was highlighted by participants rating themselves 14% more successful with Textoshop. They also found Textoshop more usable by rating it 21% higher on the SUS. This trend was reflected in participants’ comments. For example, P6 mentioned preferring Textoshop because with the baseline, “it’s what I’m using right now” and “it’s quite hard to adjust the prompts and I need […] several rounds of prompting”. P9 also mentioned that Textoshop was “easier to use and more fun to navigate” whereas with the baseline “you had to be be very, very specific otherwise it would do something completely different”. This confirms findings that people want more control when using LLMs (Masson et al., 2024; Arnold et al., 2021).
@@
Participants unanimously preferred Textoshop overall and rated it as easier for most tasks. However, four participants expressed different preferences for some tasks, justifying their choice with personal preferences. For example, P3 preferred organizing text using the baseline because “putting [text] in the margin to me, felt like using a Figma board which I don’t like”. Similarly, P7 preferred exploring variations using the baseline because “I would see [the variations] in a separate interface instead of directly modifying my text”. This points to the well-known design guideline of supporting “many paths and many styles”, which strives to offer different ways of accomplishing tasks because people work in unpredictable way (Shneiderman et al., 2006).
@@ 5.1. Observations and Participants’ Comments
5.1.1. For better or worse, some behaviours from drawing software also transferred to Textoshop
For example, participants mentioned how Textoshop allowed a trial-and-error exploration (Masson et al., 2022) similar to the exploration they were used to in drawing software. P6 mentioned “when trying to pick colours for my own drawing, I usually randomly just adjust the point on the palette and just kind of see what I can get. So I adopted the same strategy when I was trying tones”. But this sometimes also transferred difficulties. For example, regarding the boolean operations, P1 explained “Even now with photo and image editing [software], “unite” is pretty straightforward, but all the other ones […] I often can’t remember which one means what”. Finally, many participants described the system as “fun”, P3 commented “I thought it was fun just to play around with it […] it kind of sparked that same joy that I get when I mess around with image editing software”.

### v2

@@ 3.6.3. Prompting Strategies and Heuristics
Our intent with the implementation was to have robust, fast, and reliable tools to allow testing the concept behind Textoshop. While the implementation could perhaps be improved, we currently rely on engineered prompts and heuristics. For example, tools operating on a selection were optimized to only send the selection to the model and reintegrate the results by making it match the formatting of the original text. Another example is the resizing tool, which first splits the selection into sentences and prompts the model to shorten or expand each sentence in parallel. In our tests, GPT-4o was unreliable when prompted to resize a text by a certain amount of words. To boost the reliability of the resizing, the model was prompted to resize the same text eight times using a slightly different number of words. Then, the system calculated all possible combinations of the resized parts and selected the one closest to the target length.
@@ 4. User Study
We ran a two-part user study to learn about the usability and usefulness of our approach. Specifically, we tried to answer the following questions: How do people use the features of Textoshop to accomplish typical writing tasks? How does their knowledge of drawing software help them understand Textoshop? Is Textoshop easier to use than existing solutions?
@@
The second part of the study was more exploratory to observe how participants use the tool as a whole in a more holistic and unguided way. Participants used the fully-featured version of Textoshop on their own text with no other instruction than to explore the tool and its features to accomplish their text editing goals.
@@ 4.1. Participants
We recruited 12 participants from our local community (age range: 23 to 35, Mdn=26; 7 self-identified as female and 5 as male). On a 5-point scale from 1-never to 5-very frequently, participants reported how often they engaged in writing activities (Mdn=5, SD=.7), how often they use generative AI to help them during writing activities (Mdn=2.5, SD=1.5), and how often they use drawing software (Mdn=3, SD=1). Most of their writing was academic (12), professional (11), and creative (2). On a 5-point scale from 1-not at all familiar to 5-extremely familiar, they also reported their familiarity with specific drawing software features such as colour pickers (Mdn=5, SD=.8), drawing tools (Mdn=4, SD=.7), layers (Mdn=4, SD=1.1), and boolean operations on shapes (Mdn=4, SD=.6).
@@ 4.3. Procedure
Each participant used two interfaces (Textoshop and baseline). For each interface, the participant completed 4 tasks (shortening and expanding a story, organizing and integrating text fragments, changing the tone of dialogue, and using a template to write emails).
Introduction (~2min) – The participant completed the consent form and demographics questionnaire and was told they would use two different interfaces to accomplish a series of text editing tasks.
@@ 4.4. Design and Tasks
We designed four tasks such that they cover the issues with traditional ways of editing text (section 2.1) and involve as many of the features of Textoshop as possible. The tasks were designed to be realistic, require multiple steps, and ask of participants to express higher-level intents rather than simple editing operations. Rather than evaluating all tools, our main question was whether participants understand and successfully use features building on the drawing software analogy. The four tasks were as follows.
@@
Shortening/Expanding a Summary – Given a summary of a well-known novel, the participant has to shorten and expand specific sentences and paragraphs. For example, they need to compress a paragraph to avoid the short last line (“orphan”) while preserving the meaning. Or they need to expand upon a specific sentence in another paragraph so that it covers two lines and fills the whole width of the page. With Textoshop, the participant can use direct manipulation to resize text selections. This task emulates the frequent need for writers to change the length of their text so that it fits within a spatial layout (e.g., posters and presentations) or remains below a character or page limit.
@@
Changing the Tone of Dialogues – Given a dialogue between two fictional characters, the participant has to change the tone of the dialogue of one of these characters (e.g., make their dialogues more informal and positive) and then explore at least four variations of the other character’s last dialogue. With Textoshop, the participant can use the tone picker and tone-brush. This task was to test the tone picker metaphor. It emulates copyediting situations when the tone of a document has to be uniformized (e.g., integrating changes from different writers) and when the tone has to be explored precisely to match a certain intent (e.g., writing an email to different audiences).
@@
Organizing and Integrating Text Fragments – Given a simple story and a bullet list with 6 text fragments, the participant has to sort the text fragments depending on the character they are related to and then integrate these fragments into the story. Some text fragments have to be merged into specific parts of the text, while others are concepts to remove from the story. In the Textoshop condition, participants can arrange the text fragments in and off the page and can also use boolean operations to unite, intersect, subtract, or exclude text fragments. This task was to test the metaphor of the canvas and floating elements, as well as boolean operations on text fragments. It emulates editing the flow of ideas and rapidly iterating on different ways to structure these ideas.
@@
Using an Email Template to Answer Customers – Given the profile of two customers as well as the email template used by a fictional company to answer their customers, the participant has to generate an email in response to each customer. The template includes pre-written paragraphs, which should be selected if they apply to that specific customer, and has to be edited further to include the name of the customer and some specific information, such as amounts and durations. With Textoshop, layers are already created with the text of the email template; to form the final emails, the participant has to show/hide different layers and apply minor edits to the layers. With the baseline interface, the email template lists all the possible phrasings to answer a customer organized in sections; to form the final emails, the participant has to copy and paste them in order. This task was to test the layer metaphor and participants’ mental model of layers when editing. It simulates having to produce different text versions simultaneously based on different audiences, such as email communication or educational resources.
@@ 4.5. Data Analysis
Considering ongoing debates about statistics in HCI (Dragicevic, 2016; Group, 2019), we report both confidence intervals (CI) and p-values. However, we base our conclusions on the 95% confidence intervals (CI) on mean differences, which provide information about both the presence and size of effects. All confidence intervals are calculated using the studentized bootstrapping method because it is the most robust for within-subject study designs of about 12 participants (Masson et al., 2023; Zhu and Kolassa, 2018). All p-values are calculated using a Wilcoxon signed-rank test because it is less likely to yield false positives compared to parametric alternatives when dealing with data whose distribution is unknown and possibly heavily skewed (Bridge and Sawilowsky, 1999). The bootstrapped CIs are calculated in Python using arch 5.5.0 (Sheppard et al., 2023) and using 10,000 replications. The Wilcoxon signed-rank tests are calculated using scipy 1.10.1 (Virtanen et al., 2020) with ties being discarded.
@@ 4.6. Results from the Text Editing Tasks
Stacked bar charts of the responses from participants to the 5-point scale statements. The majority of the ratings for Textoshop are in the agree and strongly agree zone, whereas the ratings for the baseline are more split, with some questions such as ”Precisely shortening/expanding text was easy” being mostly rated strongly disagree and disagree.
@@ 4.6.2. Usability
4.6.3. Time and Prompts Used
Although productivity was not our focus, we provide indications of completion time and prompts used as a loose indicator of the tedium involved in completing a task. Participants completed tasks faster using Textoshop than baseline by 55s (95% CI [26s, 1min 26s], M=2min 23s vs M=3min 19s, p=0.003). Ranked from largest to smallest effects: “Shortening/expanding a summary” by 1min 29s, “Using an email template” by 1min 19s, “Changing the tone of dialogues” by 41s, and “Organizing and integrating text fragments” by 13s. Additionally, in the baseline condition, participants used 2.7 prompts per task on average, and these prompts were about 269 characters long. Considering participants did not use prompts in the Textoshop condition, we discuss participants’ behaviours in section 5.1.
@@ 4.6.5. Preferences
When asked about their preferred interface for each task, participants unanimously rated Textoshop as “preferred” or “strongly preferred” when “integrating text fragments”. All participants also preferred Textoshop when “precisely shortening/expanding text” except P10, who did not have a preference for either interface. For other tasks, while the majority preferred Textoshop, some participants expressed other preferences: when “writing using a template” P8 preferred the baseline interface; when “organizing text fragments” P3 preferred and P8 strongly preferred the baseline; and when “exploring meaningful variations of text” P7, P8, and P12 preferred and P3 strongly preferred the baseline. We discuss the participants’ reasons for these preferences in Section 5.
@@ 4.7. Results from the Free-form Exploration
In the free-form exploration, participants chose texts related to creative story writing (P2, P4, P5, P9), informative writing (P3, P10, P11, P12), academic writing (P1, P6, P7), and personal writing (P8). Figure 7 shows the sequence of features used by each participant. Below, we report some of the strategies and behaviours observed.
Figure 7. Sequence of operations done by participants using Textoshop during the free-form exploration. Repeated sequences of the same feature are represented as a stacked pile. Refer to section 3 for the meaning of the icons.
[…]
List of commands used by each participant during the free-form part of the study. Overall, participants covered all types of commands. Boolean operations were the least frequent, with P3, P4, P7, P8, P9, and P10 never using them. P4 and P10 are the only ones that did not use layers. And P8 was the only one that did not use the tone wheel. Direct manipulations were used by all at least once, if only to move or resize sentences.
@@ 4.7.1. Iterating on a sentence by alternating tools
Participants used the tools to iterate on the wording of some specific sentences, much like they would iterate on the colours, texture, and organization of elements in drawing software. However, unlike visual artists who often apply colours and texture only after sketching (Tsandilas et al., 2015), with Textoshop, tone exploration happened right from the beginning. Often, participants’ workflow involved alternating between trying different tones, shortening the sentence, and smudging some parts to be paraphrased. The order of these steps was not always the same, but the tone picker was often used first, and then the other tools would allow editing with more precision. With the exception of P2, who frequently rotated and split the sentences, other tools were seldom used to rewrite a sentence.
@@ 4.7.2. Experimenting with alternative organizations of ideas
P5, P10, and P12 reorganized the ideas multiple times by taking a passage from one paragraph and integrating it into a different paragraph or sentence using the boolean operations. P6 did the same but created their own text fragments to add an idea to an existing paragraph. Finally, P1 went one step further and deconstructed a whole paragraph into passages. Then, they organized these passages in the margin to try a different organization of the arguments and then reintegrated the passages into the main text using the “insert” and “unite” features. This suggests Textoshop encouraged writers to work on the flow of ideas even after the writing is fully-fledged into sentences and paragraphs. This is in contrast with traditional writing, where restructuring mostly happens in the planning phase, when the draft is only an outline (Flower and Hayes, 1977).
@@ 4.7.3. Using layers to identify changes, store AI modifications, or add structure
Layers were used in multiple ways, most differing from how they are used in drawing software. For example, P1, P8, and P9 used one layer to store all the modifications made using the tools, allowing them to preserve the integrity of their original text and easily identify AI changes by switching the visibility of the layer. P3 used layers to store alternative versions of paragraphs (in this case, storing a more informal version). Perhaps the closest use of text layers to drawing layers that we have observed is to use them to structure the text. For example, P5, P11, and P12 created layers to store the different paragraphs, passages, or main ideas. P7 used layers to store the examples mentioned in the text. And P6 used a layer to store the statistics reported in the text so that they could read and edit their results section while hiding the means or p-values reported in the text (by hiding the statistics layer). This differs from traditional writing, where text cannot usually be teased apart in semantically meaningful ways. It is akin to the organization proposed by systems like Scrivener (Scrivener, 2005) except with a much finer granularity than just chapters and sections.
4.7.4. Non-linear editing and reviewing
While it is known that reviewing and editing are non-linear processes (Flower and Hayes, 1977), Textoshop seemed to be more favourable to this behaviour than existing word processors. P3, P4, P6, P7, P9, P11, and P12 seemed to follow a tool-driven approach to editing, where selecting a new tool triggered a new pass on the document. For example, after attending to changes at the end of the document, P1 selected the eraser and went through the document from the beginning, this time trying to erase words that may not be essential, such as adverbs. Still, P2, P5, P8, and P10 proceeded more linearly, reading through the documents and switching between tools as needed, in a process akin to editing in a traditional word processor.
@@ 5. Discussion
We first answer our main research questions using the results and participants’ comments. Then, we open the discussion to themes stemming from observations and discussions with participants.
Participants successfully used Textoshop to accomplish writing tasks in ways that differ from traditional writing
This was highlighted by participants rating themselves 14% more successful with Textoshop. They also found Textoshop more usable by rating it 21% higher on the SUS. This trend was reflected in participants’ comments. For example, P6 mentioned preferring Textoshop because with the baseline, “it’s what I’m using right now” and “it’s quite hard to adjust the prompts and I need […] several rounds of prompting”. P9 also mentioned that Textoshop was “easier to use and more fun to navigate” whereas with the baseline “you had to be very, very specific otherwise it would do something completely different”. This confirms findings that people want more control when using LLMs (Masson et al., 2024; Arnold et al., 2021).
@@
Participants unanimously preferred Textoshop overall and rated it as easier for most tasks. However, four participants expressed different preferences for some tasks, justifying their choice with personal preferences. For example, P3 preferred organizing text using the baseline because “putting [text] in the margin to me, felt like using a Figma board which I don’t like”. Similarly, P7 preferred exploring variations using the baseline because “I would see [the variations] in a separate interface instead of directly modifying my text”. This points to the well-known design guideline of supporting “many paths and many styles”, which strives to offer different ways of accomplishing tasks because people work in unpredictable ways (Shneiderman et al., 2006).
@@ 5.1. Observations and Participants’ Comments
5.1.1. The drawing analogy was mostly successful
Good interface analogies “help establish user expectations and encourage predictions” (Neale and Carroll, 1997). We observed such expectations and predictions from our participants, as they mentioned how Textoshop allowed a trial-and-error exploration (Masson et al., 2022) similar to the exploration they were used to in drawing software. P6 mentioned “when trying to pick colours for my own drawing, I usually randomly just adjust the point on the palette and just kind of see what I can get. So I adopted the same strategy when I was trying tones” suggesting part of the success of the colour picker analogy is that it matched how one would use it in drawing software. However, one danger with interface analogies is the “overly literal translation of existing bad designs” (Preece et al., 2002). We observed one instance of such phenomenon regarding the boolean operations, P1 explained “Even now with photo and image editing [software], “unite” is pretty straightforward, but all the other ones […] I often can’t remember which one means what”. Finally, many participants described the system as “fun” and P3 commented “I thought it was fun just to play around with it […] it kind of sparked that same joy that I get when I mess around with image editing software”. This relates to the observation from Carroll and Thomas (1982) that an interface analogy that “is conducive to the desired emotional attitude of the user” and “provide the user with exciting metaphors for routine work” may enhance the intrinsic motivation of users in accomplishing their tasks.
5.1.2. Participants quickly learned how to use Textoshop, but it is unclear if they know when and how to integrate it into their workflow.
The user study tested the first part of the learning curve of Textoshop, which corresponds to the time it takes for a user to reach a reasonable level (Nielsen, 1994). Our results and observations suggest participants started high on the learning curve, possibly thanks to transferring their knowledge from drawing software (Polson et al., 1986; Neale and Carroll, 1997). However, we only have anecdotal evidence about the shape of the second part of the learning curve, which corresponds to the transition from novice to expert. Specifically, we observed participants had no issues trying features and workflows and undoing them if necessary, which might help discover new features and expert workflows (Masson et al., 2022). However, it remains unclear if people would actually integrate features of Textoshop into their workflow. For example, it is known that, despite knowing more efficient ways of accomplishing tasks, such as keyboard shortcuts, users might still prefer their inferior approach (Cockburn et al., 2015; Banovic et al., 2012). Similarly, users and Textoshop might co-adapt such that features of Textoshop are appropriated and used in unexpected ways (Mackay, 2000).